OpenStudy (anonymous):

A teacher sells children's tickets for $4 and adult tickets for $7 . One night 575 tickets worth $3575 were sold. How many adults and how many childrens bought tickets?

OpenStudy (anonymous):

x = 9/8 and y = -3/2